OpenStudy (anonymous):

what is the square root of (4x^6) simplified?

OpenStudy (e.mccormick):

\[\sqrt{4x^6}\]?

geerky42 (geerky42):

4 = 2², \(\large x^6 = (x^3)^2\) so \(\sqrt{4x^6} = \boxed{|2x^3|}\)

geerky42 (geerky42):

Is this clear? @BridgetF

OpenStudy (anonymous):

sorry, I still dont understand

OpenStudy (e.mccormick):

OK. What he has done is change 4 into 2 squared and the x to the 6th to x to the 3rd that is squared. Do you know the power rules for doing that? Or is the confusion elsewhere?

OpenStudy (anonymous):

The confusion is everywhere. I dont seem to get any of this. I am exhausted from thinking and not understanding

OpenStudy (e.mccormick):

That's OK. It happens. By the power rules, a power times a power is the powers added and a power to a power is the powers multipled. For example: \[a^2\times a^3 = a^{2+3} = a^5 \;and\; (b^3)^5=b^{3 \times 5}=b^{15}\] Geerky42 was using those rules to play around with things. That made the 4 into 2 squared and that other thing he did: \[(x^3)^2\]

OpenStudy (anonymous):

why is it an absolute value at the end?

OpenStudy (e.mccormick):

Because of the square root and powers of 2 Any power of 2 will make something positive. so (-1)^2 is 1 and (-2)^2 is 4 and so on. Because of that, the answer would always be positive.

OpenStudy (anonymous):

Thanks for your help :-)

OpenStudy (e.mccormick):

np. I hope that all makes more sense now. If not, you may want to look into the atributes of exponents. Here is a site that has all sorts of stuff like this written out a little different than most books: http://www.purplemath.com/modules/exponent.htm I find the second way of talking about it helps me.